OpenStudy (anonymous):

One professor grades homework by randomly choosing 5 out of 12 homework problems to grade. --- Jerry did only 5 problems of one assignment. What is the probability that the problems he did comprised the group that was selected to be graded?

OpenStudy (anonymous):

pretty damned small

OpenStudy (anonymous):

\[\frac{5}{12}\times \frac{4}{11}\times \frac{3}{10}\times \frac{2}{9}\times \frac{1}{8}\]

OpenStudy (anonymous):

\[=\frac{1}{11\times 9\times 8}\] if my arithmetic is right